# LLM-Agent-UMF: LLM-based Agent Unified Modeling Framework for Seamless Integration of Multi Active/Passive Core-Agents

### v1

@@ 4.1 Evaluation of the new core-agent terminology
Agents have traditionally been discussed as a single monolithic unit, with different inner functional entities, such as LLMs and complementary software components, treated in an intertwined manner, lacking clear boundaries definition. This approach can lead to confusion when attempting to analyze or compare various systems that exhibit agent-like behaviors. One of the main reasons why agent definition remains misunderstood lies in the disparity between how researchers and practitioners use terminology to describe their work. For instance, some studies might not explicitly refer to their LLM-based system as an "agent", even though it exhibits characteristics commonly associated with agents such as autonomy, adaptability, and goal-directed behavior. The ToolLLM [54] research paper provides a prime example of this discrepancy. While the authors indeed utilize API retriever component in conjunction with their LLM, they refrain from using the term "agent." The same can be observed in the Toolformer [37] work, which implicitly incorporates software to execute API call requests generated by LLMs but similarly avoids referencing the system as an agent.
@@
One terminological ambiguity we resolve in this work pertains to the concept of “memory”. For instance, ChatDB [30] illustrates a dual usage by defining memory in one context as the data retrieved by the agent from knowledge sources, while in other contexts, it refers to the space for storing processed data and information essential to the core-agent’s operations. To distinguish clearly between these usages, we clarify that the former definition pertains to a "data repository," whereas the latter accurately corresponds to "memory."
In addition to resolving ambiguities, based on the definition that we established for the “core-agent” component, we successfully identified core-agents within multiple LLM-based systems, mainly Toolformer and ToolLLM. On the one hand, Toolformer incorporates a simple-structured core-agent that includes only the action module responsible for API execution.
On the other hand, ToolLLM adopted a more sophisticated approach. In fact, we recognize the use of a core-agent encompassing two modules: an action module represented by the neural API retriever and an implicate planning module responsible of managing the flow of information within the agent: The core-agent intercept user instructions, leverage the API retriever to gather relevant APIs, relays the APIs to the LLM for response generation, executes the requested APIs, and finally returns the outcome back to the LLM for final user response formulation. Thus, the identification of core-agents in these systems, accompanied with the presence of LLMs elucidates that Toolformer and ToolLLM are indeed LLM-based agents.
@@
The integration of the security module into the core-agent is substantiated by the presence of security concerns in LLMs and the imperative to address them comprehensively. While some studies focus solely on guardrail techniques or algorithms without tying them to an agent, the application of guardrails leveraging our framework occurs within the frame of a core-agent, particularly within its security module.
@@
For example, LLMSafeGuard [49] introduces a lightweight framework to protect in real-time the LLM text generation. This study opted for the usage of the beam search algorithm to generate candidate responses. Leveraging an external validator, it rejects candidates that violate security constraints and proceeds with valid ones. Projecting this work on our framework, we concluded that the external validator is indeed a core-agent whose planning module assesses the alignment of a candidate with the agent’s security constraints to determine whether to generate alternative candidates in case of rejection or proceed with sentence completion in case of acceptance. Evidently, this module communicates with the security module which implements a real-time validation approach to decide which path to take in the LLM’s generation process.
@@
Despite paper [60] emphasizes on the necessity of achieving reliability, confidentiality, and integrity in LLM-based agents by proposing methods to attain each attribute, it does not explicitly discuss the implementation of these mechanisms architecturally. Additionally, all discussed techniques serve the diverse security goals outlined in our framework, reinforcing the rationale for relocating guardrails to the security module within the core-agent.
@@
4.2 Evaluation of active/passive core-agent concept
@@
Our introduced framework, LLM-Agent-UMF, provides a valuable tool for comparing several state-of-the-art LLM-based agents as represented in Table 3. In this section, we evaluate our LLM-Agent-UMF by projecting it onto existing agent implementations and highlighting the presence or absence of essential modules such as planning, memory, profile, action, and security modules. This analysis allows us to classify the identified core-agents into active and passive. Besides, it provides insights into existing agents structures facilitating opportunities to merge functionalities from different agents within one system for enhanced capabilities.
Since some solutions such as ChatGPT are proprietary and closed-source, we cannot analyze its structure and capabilities assertively or project them transparently onto our own framework. Based on the observed behavior and public information about ChatGPT, we can make plausible hypotheses regarding its features and capabilities, but we cannot state them definitively.
To guide our hypothesis, we assessed ChatGPT 4o mini’s response to illegal prompts by demanding the steps to hotwire a car. Our investigation aimed to verify the system’s behavior when confronted with malicious queries. Matching our expectation, ChatGPT 4o mini declined to respond to direct illegal prompts, as illustrated in Appendix A, Figure 14. Consequently, we formulated two hypotheses on how the security measures are implemented:
  • Hypothesis 1: The LLM itself is trained to monitor the input and prevent generation of undesirable output. This can be achieved by adhering to Adversarial Training (AT) techniques [61] and implementing input validation and sanitization to prevent malicious prompts from being used. Furthermore, if we consider ChatGPT 4o mini as an agent, there is no clear evidence of its interaction with external tools, which would necessitate functionalities of an action module in a core-agent. Based on this argument, we hypothesize that ChatGPT 4o mini is not an agent in the first place.
Similarly, we consider the capability of ChatGPT 4o on code execution and suppose the presence of an active core-agent. Indeed, we constructed the third hypothesis, around that process as follows: the core-agent checks if the prompt requires coding operations. If it is the case, it changes the profile of the LLM according to the task at hand, then leverages the LLM capability to generate code which is later executed in an isolated environment. Afterwards it reset the profile of the LLM to explain the results in a suitable textual representation. This hypothesis could be further enriched with assumptions from the ChatGPT 4o mini second hypothesis about the security measures and leads us to the conclusion that ChatGPT 4o is an LLM-based agent powered with a fully featured active core-agent rather than a standalone LLM.
In table 3, we categorize a group of LLM-based agents. Our selection includes diverse open-source and proprietary agents.
Table 3: Classification of state-of-the-art agents using LLM-Agent-UMF
@@
Planning Profile Memory Action Security Core-Agent Category
ToolLLM [54] X [*] X X X - Active
Zero-shot - - - X - Passive
Gorilla [62] [**] With retriever X [*] X X X - Active
@@
GPT4Tools [65] [***] X X X X - Active
Chameleon [7] X X X X - Active
ChatDB [30] X X X X - Active
LLM+P [59] X X X X - Active
LLMSafeGuard [49] X - X X X Active
Hypothesis 1 [****] - - - - - N/A
ChatGPT 4o mini Hypothesis 2 [****] X - - X X Active
ChatGPT 4o Hypothesis 3 [****] X X X X X Active
[∗] Both ToolLLM and Gorilla (with retriever mode) use a retriever that recommends APIs and fetch their documentation.
[∗∗] Gorilla was implemented in 2 versions: the first version is “zero-shot” without prompt tuning, the second is with a retriever which retrieves up-to-date API documentation to relay back to the LLM.
[∗∗∗] GPT4Tools framework leverages multi-modal tool usage. In case of a multi-step subtask, any procedure resulting in an intermediary output is stored by the memory module and employed by the planning module to reach the final output.
[∗∗∗∗] ChatGPT is a proprietary software owned by OpenAI, so we have formulated hypotheses regarding its structure, as detailed in the paragraph preceding Table 3.
@@
By examining the core-agents within each agent independently, developers can deduce the compatibility of core-agents and identify challenges that may arise when integrating multiple agents into one system such as the necessity of synchronization, potential conflicts between core-agents or functional redundancy.
@@ 4.3 Evaluation of multi-core agent architectures
Thanks to the classification done in table 3, we can now effortlessly merge various aspects from existing agents into a single entity. To demonstrate this capability, we propose four representative scenarios that highlight the potential of LLM-Agent-UMF for designing multi core-agent systems by combining distinctive features from state-of-the-art agents. Each of these scenarios utilizes Llama 3.1 8B [53], the newest state-of-the-art LLM from Meta AI team, for its remarkable performance and optimized memory footprint.
4.3.1 Toolformer and Confucius as a multi passive core-agent system
@@ 4.3.2 ToolLLM and ChatDB as a multi active core-agent system
4.3.3 Implanting LLMSafeGuard security module in ToolLLM
@@
4.3.4 Hybrid active/passive core-agent system
@@
The four previous scenarios exemplify how researchers and developers can make informed decisions about the design of their LLM-based agents prior to the development process, grounded in clear architectural reasoning. This systematic approach will enhance the robustness and functionality of LLM-based agents. In the next section, we will discuss the limitations and the future work to enhance our framework.
Figure 13: LLM-based Agent 4 (LA4): LLM+P, LLMSafeGuard, Toolformer and Confucius –
Hybrid Active/Passive Core-Agent

### v2

@@ 4 Results and discussion
LLM-based agents have traditionally been discussed and conceptualized as intricate systems with different inner functional entities, such as LLMs and complementary software components, treated in an intertwined manner. Our proposed LLM-based Agent Unified Modeling Framework (LLM-Agent-UMF) was shaped to overcome these issues and promote a clear delineation of components and responsibilities. In this section, we present and analyze the outcomes derived from adopting the LLM-Agent-UMF through diverse evaluation approaches. Firstly, Section 4.1 evaluates the newly introduced core-agent terminology by applying the LLM-Agent-UMF to existing systems that do not explicitly identify themselves as agents. Secondly, Section 4.2 dives deeper into understanding and dissecting the internal structure of thirteen LLM-based agents by leveraging our proposed framework. This is performed not only on open-source agents but also on proprietary ones by formulating hypothesis based on their observed behavior. Through this process, we classify core-agents into active and passive types and identify critical modules that may have been overlooked or underemphasized during development. Finally, Section 4.3 presents five novel architectures based on the LLM-Agent-UMF to model multi-core agents by combining characteristics from separate agents not intended for fusion in their original design.
@@ 4.1 Evaluation of the new core-agent terminology
One of the main reasons why agent definition remains misunderstood lies in the disparity between how researchers and practitioners use terminology to describe their work. For instance, some studies might not explicitly refer to their LLM-based system as an "agent", even though it exhibits characteristics commonly associated with agents such as autonomy, adaptability, and goal-directed behavior. The ToolLLM [59] research paper provides a prime example of this discrepancy. While the authors indeed utilize API retriever component in conjunction with their LLM, they refrain from using the term "agent". The same can be observed in the Toolformer [38] work, which implicitly incorporates software components to execute API call requests generated by LLMs but similarly avoids referencing the system as an agent.
@@
Based on the definition that we established for the core-agent component, we successfully identified core-agents within multiple LLM-based systems, mainly Toolformer and ToolLLM. On the one hand, Toolformer incorporates a simple-structured core-agent that includes only the action module responsible for API execution. On the other hand, ToolLLM adopted a more sophisticated approach. In fact, we recognize the use of a core-agent encompassing two modules: an action module represented by the neural API retriever and an implicate planning module responsible of managing the flow of information within the agent: The core-agent intercept user instructions, leverage the API retriever to gather relevant APIs, relays the APIs to the LLM for response generation, executes the requested APIs, and finally returns the outcome back to the LLM for final user response formulation. Thus, the identification of core-agents in these systems, accompanied with the presence of LLMs elucidates that Toolformer and ToolLLM are indeed LLM-based agents.
@@
It’s worth noting that another paper introduced a framework called ChatDB [31] which leverages databases as symbolic memory for LLMs. Although they referenced other memory-augmented LLM techniques such as Auto-GPT [65] and Generative Agent [66], the authors did not categorize their own framework as an agent. They presented a system comprising two components: an LLM controller and a memory module. Despite mentioning that any commonly used LLM can be employed as the controller component, it is evident that the LLM alone cannot interface with its associated database memory extension without the need of an intermediary software element. This analysis prompted us to identify an inherent core-agent responsible for managing the integration of LLMs into SQL query generation processes and executing these queries in the context of databases representing the system’s memory. Following this analysis, we deduce that ChatDB aligns with our defined software architectural framework for an LLM-based agent.
@@
The integration of the security module into the core-agent is substantiated by the presence of security concerns in LLMs and the imperative to address them comprehensively. While some studies focus solely on guardrail techniques or algorithms without tying them to an agent, by leveraging our framework, the application of guardrails occurs within the frame of a core-agent, particularly within its security module.
@@
For example, LLMSafeGuard [54] introduces a lightweight framework to protect in real-time the LLM text generation. This study opted for the usage of the beam search algorithm to generate candidate responses and leverage an external validator to handle the safety checking. Projecting this work onto our framework enabled us to conclude that this study indeed describes an LLM-based agent orchestrated by an inherent core-agent. The core-agent utilizes the LLM to generate candidate response, assesses its alignment with the agent’s safety constraints and evaluates whether to proceed with sentence completion in case of acceptance or generate alternative candidates in case of rejection.
@@
Besides, despite paper [67] emphasizing on the necessity of achieving reliability, confidentiality, and integrity in LLM-based agents, it does not explicitly discuss the implementation of these mechanisms architecturally. Additionally, all discussed techniques serve the diverse security goals outlined in our framework, reinforcing the rationale for relocating guardrails to the security module within the core-agent.
The successful identification of core-agents within systems that do not identify themselves as agents emphasizes the significance of establishing a consistent terminology to accurately describe entities and underscore the importance of assigning distinct roles to software components within complex systems.
@@
4.2 Evaluation of active/passive core-agent internal structure delineation
@@
Our introduced framework, LLM-Agent-UMF, constitutes a valuable tool for comparing several state-of-the-art LLM-based agents as represented in Table 3. In this section, we evaluate our LLM-Agent-UMF by projecting existing agents onto our framework, distinguishing core-agents and highlighting the presence or absence of essential modules such as planning, memory, profile, action, and security modules.
Table 3: Classification of state-of-the-art agents using the LLM-Agent-UMF
@@
Planning Profile Memory Action [*] Security Core-Agent Category
@@
Zero-shot - - - X - Passive
Gorilla [70] With retriever X - M X - Active
ToolLLM [59] X - M X - Active
GPT4Tools [36] M - M X - Active
Chameleon [8] X X M X - Active
ChatDB [31] X X X M - Active
ChemCrow [12] X X M X M Active
LLM+P [64] X X M M - Active
LLMSafeGuard [54] X - - M X Active
Hypothesis 1 - - - - - N/A
ChatGPT 4o mini Hypothesis 2 X - - M X Active
ChatGPT 4o Hypothesis 3 X X X X X Active
X: Denotes the presence of a module and the fact that its functionalities were well discussed in the research.
M: Denotes the presence of a minimal implied module and that it was not the main focus of the research.
[∗] The action module is an essential module in a core-agent. In its minimal form, it is responsible for human-machine interaction only.
In this case study, the thirteen agents that will be thoroughly examined are Toolformer [38], Confucius [68], ToolAlpaca [69], Gorilla [70], ToolLLM [59], GTP4Tools [36], ChatDB [31], Chameleon [8], LLM+P [64], ChemCrow [12], LLMSafeGuard [54], as well as both ChatGPT 4o and its minimal version. Each of these LLM-based agents has been carefully selected to showcase a diverse range of functionalities across various modules within our proposed framework. The outcome of this dissection constitutes a valuable medium to identify overlooked functionalities and complementary counterparts. It also allows us to distinguish entities that may either play or not an authoritative role in a system, subsequently categorizing them as active or passive core-agents.
Initially, our analysis focuses on Toolformer [38], Confucius [68] and ToolAlpaca [69] which possess significant similarities in terms of functionality and behavior. All three agents primarily offer innovative finetuning methods aimed at influencing LLMs to utilize APIs for more accurate results. Notably, they do not require external assistance from other software components for orchestration or API selection; instead, the LLM’s inherent capabilities are leveraged for this purpose. The passive core-agent design within these tools is notable, featuring a unique action module responsible for executing the LLM’s will in making API calls and relaying responses back to the model for improved formulation.
Going further in our analysis, we study ToolLLM [59] and Gorilla [70], and we examine the latter two modes: "zero-shot" and "with retriever". In the zero-shot mode, Gorilla behaves similarly to the three previously discussed agents by acting as an LLM-based agent with a passive core-agent that possesses an action module for API call execution. However, in the second mode, functioning similarly to ToolLLM, Gorilla utilizes an API retriever to gather API recommendations and fetch documentation. This process necessitates a software component for orchestrating interactions with LLMs and tools; this component, which we label as the core-agent, was not highlighted in the original papers. It is inherent that this core-agent includes a planning module incorporated as the API retriever and requires a minimal memory module to manage the internal state of the agent in case of multi-step instructions.
We complement our analysis by deeply evaluating the case of GPT4Tools [36], which took inspiration from its predecessors and focused on finetuning the LLM to utilize tools but extended its capabilities to leverage visual models such as image segmentation and generation models. Once trained using their technique, the LLM is capable of responding to instruction by performing multiple steps to achieve a targeted goal. To satisfy the needs of such an LLM, a software component must be integrated to listen to its requests, execute actions and either return textural responses or store intermediate outputs such as images in its memory to be utilized in subsequent steps. This software entity, identified as the core-agent, lacks any established authoritative nature but requires basic planning and memory modules to manage multi-step subtasks. Therefore, we classify it as an active core-agent.
Projecting ChatDB [31] onto our framework, as discussed in Section 4.1, reveals the existence of a core-agent. Its managerial position within the agent system makes a clear indication of its active nature. In fact, this core-agent incorporates a planning module that follows a chain-of-memory approach and utilizes an LLM for decomposing complex problems into multiple steps of memory operations. To efficiently manage read and write operations to the memory extension implemented as a SQL database, an additional memory module is essential. Moreover, we highlight the necessity of a profile module adopting the in-context learning approach to influence the LLM in generating appropriate SQL queries.
Additionally, Chameleon [8], another LLM-based agent, harnesses the power of GPT-4 for planning and selecting suitable tools to achieve desired goals. It synthesizes programs that leverage various tools such as LLMs, vision models, python functions or heuristic-based modules. To accomplish this, a software component is necessary to manage the flow of information and utilize these tools effectively. In this context, we identify a core-agent comprising: a planning module empowered with GPT-4, a profile module to control the behavior of the LLM without any prior training, a minimal memory module to store intermediate results and a versatile action module that make use of the available tools. The internal structure and the managerial aspect of this core-agent lead us to classify it as an active core-agent.
As previously elucidated in Section 4.1, the LLM+P [64] solution is recognized as an LLM-based agent processing a core-agent with a sophisticated planning module integrating a PDDL interpreter. The described approach requires influencing GPT-4 to convert from and to PDDL representation and thus underscore the need for profile module. The core-agent needs also a minimal memory module to manage the internal state of the agent as there are multiple phases in the planning process. The presence of these modules emphasizes the active nature of the core-agent.
Furthermore, we would like to emphasize the significant effort invested in ChemCrow [12], a chemistry-oriented agent. When projecting it onto LLM-Agent-UMF, we recognize ChemCrow as an active core-agent itself. The researchers explicitly defined it as an independent entity from both the LLM and the available tools. Notably, this intricate core-agent leverages the capabilities of GPT-4 to coordinate the utilization of 18 expert-designed tools for synthesizing chemical compounds. As a result, we acknowledge the presence of several crucial modules within ChemCrow: A planning module responsible for orchestrating the entire process; a minimal memory module that stores intermediate results; a profile module to guide GPT-4’s text generation process; a versatile action module managing communication with all available tools; and a security module, which despite its simplicity, plays a vital role in checking safety information before proceeding further with the task at hand and interrupts the execution if it is deemed dangerous.
Another notable solution is LLMSafeGuard [54] that we have already identified as an LLM-based agent with an implicit core-agent. The researchers augmented the beam search algorithm with an external validator which rejects candidates that violate security constraints and proceeds with valid ones. The core-agent in this solution is indeed an active one and includes a planning module, which ensures the correct execution of the workflow, and a security module represented by the external validator.
As a prominent language modeling solution, ChatGPT’s inner workings warrant thorough examination to better understand its capabilities and limitations. However, being a proprietary and closed-source system limits our ability to assertively analyze its structure and capabilities. Although it exhibits characteristics that align with a simple LLM, we cannot definitively conclude whether it is indeed a standalone LLM or an LLM-based agent. Nevertheless, based on public information and rigorous analysis of ChatGPT behavior through the lens of our framework, we can make plausible hypotheses regarding its structure.
Figure 14: ChatGPT 4o mini refusing to explain how to hot-wire a car
To guide our hypothesis, we assessed ChatGPT 4o mini’s response to illegal prompts by demanding the steps to hotwire a car. Our investigation aimed to verify the system’s behavior when confronted with malicious queries. Matching our expectation, ChatGPT 4o mini declined to respond to direct illegal prompts, as illustrated in Figure 14. Consequently, we formulated two hypotheses on how the security measures are implemented:
  • Hypothesis 1: The LLM itself is trained to monitor the input and prevent generation of undesirable output. This can be achieved by adhering to Adversarial Training (AT) techniques [71]. Additionally, it has been verified that ChatGPT 4o mini does not possess direct access to external tools or sources of information; hence, we can hypothesize that it is not an agent in the first place.
Similarly, we consider the capability of ChatGPT 4o on code execution and suppose the presence of an active core-agent. Indeed, we construct the third hypothesis, around that process as follows: the core-agent checks if the prompt requires coding operations. If it is the case, it changes the profile of the LLM according to the task at hand, then leverages the LLM capability to generate code which is later executed in an isolated environment. Afterwards it reset the profile of the LLM to explain the results in a suitable textual representation. This hypothesis could be further enriched with assumptions from the ChatGPT 4o mini second hypothesis about the security measures and leads us to the conclusion that ChatGPT 4o is an LLM-based agent powered with a fully featured active core-agent rather than a standalone LLM.
Through this exercise, we demonstrate how the LLM-Agent-UMF can assist developers in reevaluating their agent designs and potentially improving upon them. Specifically, out of all studied agents that utilize external tools, 78% (7 over 9) did not incorporate necessary security measures to handle privacy concerns. Only ChemRow and ChatGPT demonstrated an ability to manage this aspect effectively. ChemRow utilized IBM Research’s tools, which fundamentally emphasize privacy protection [72], while ChatGPT takes an additional step by requiring user validation before sharing any data with third-parties. This highlights the critical importance of addressing privacy concerns and safeguarding against information leakage to ensure robustness and trustworthiness in AI-powered systems.
In our analysis, we identified that only 31% (4 out of 13) of the examined agents utilized passive core-agents. Conversely, a majority of 69% had active core-agents. This distinction may be significant when considering integrating these agent types into multi-core systems as passive core-agents are not typically subject to heavy synchronization processes, which can potentially improve system scalability.
Moreover, we observed that while some modules were present in the studied agents, they were sometimes implemented minimally and could likely be easily merged with other systems offering more comprehensive module implementations. For instance, ToolLLM has a simplistic memory module compared to ChatDB’s robust version. It is also crucial to note that the planning module plays an integral role in both agents. Therefore, any merge or integration must consider these distinct features to maximize synergy and prevent compromising functionality within the resulting system.
@@
By examining the core-agents within each agent, researchers and practitioners can deduce the compatibility of core-agents and identify challenges that may arise when integrating multiple agents into one system such as the necessity of synchronization, potential conflicts between core-agents or functional redundancy.
@@ 4.3 Evaluation of multi-core agent architectures
Thanks to the classification done in Table 3, we can now effortlessly merge various aspects from existing agents into a single entity. To demonstrate this capability, we propose five representative scenarios that highlight the potential of LLM-Agent-UMF for designing multi core-agent systems by combining distinctive features from state-of-the-art agents. Each of these scenarios utilizes Llama 3.1 8B [58], the newest state-of-the-art LLM from Meta AI team, for its remarkable performance and optimized memory footprint.
4.3.1 Toolformer and Confucius as a multi passive core-agents system
@@ 4.3.2 ToolLLM and ChatDB as a multi active core-agent system
Figure 17: LLM-based Agent 2-B (LA2-B): ToolLLM and ChatDB – Monolithic active core-agent architecture
4.3.3 Implanting the LLMSafeGuard security module into ToolLLM
@@
4.3.4 Hybrid multi active/passive core-agents system
@@
Figure 19: LLM-based Agent 4 (LA4): LLM+P, LLMSafeGuard, Toolformer and Confucius - Hybrid multi active/passive core-agents architecture
@@
The five discussed scenarios exemplify how researchers and developers can make informed decisions about the design of their LLM-based agents prior to the development process, grounded in clear architectural reasoning. This systematic approach will enhance the robustness and functionality of LLM-based agents. In the next section, we will discuss the limitations and the future work to enhance our framework.